OpenStudy (anonymous):

Solve: −4(x + 1) − 3 = −3(x − 4)

OpenStudy (anonymous):

-4x-4-3=-3x+12 -4x-7=-3x+12 -19=x or x=-19

OpenStudy (anonymous):

mul -4 with inner terms n -3 with inner terms bring x terms one side n constacts one sides n proceed

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

I love seeing such helpful people!